OpenStudy (anonymous):

Holy moly! How do I do this one?? (Please give me a moment to insert the equation)

OpenStudy (anonymous):

\[\LARGE 4\sqrt[4]{5} + 7\sqrt[3]{5} - 9\sqrt[3]{5} + 6\sqrt[4]{5}\]

hartnn (hartnn):

just combine like terms...this is simpler

OpenStudy (anonymous):

\[\large A) 8\sqrt{5}\] \[\large B) 8\sqrt{10}\] \[\large C) 10\sqrt[4]{10} - 2\sqrt[3]{10}\] \[\large D) 10\sqrt[4]{5} - 2\sqrt[3]{5}\]

OpenStudy (cwrw238):

add like terms eg